# InSyn: Modeling Complex Interactions for Pedestrian Trajectory Prediction

### v1

@@ 4.3 SSOS: Seq-Start of Seq
4.4 Seq-CVAE: LSTM-Based CVAE
RNN (typically LSTM [11]) and Conditional Variational Autoencoder (CVAE) [26] have been widely applied in pedestrian trajectory prediction [15, 40, 43, 1]. We combine the two approaches and propose a generator, Seq-CVAE (Sequence CVAE), which is designed for sampling trajectory goals. The model predicts goals given the condition $c$, which is derived from the observed trajectory $pos_{0:\tau}$. These goals are subsequently fed into the Interaction Encoder to achieve its goal-driven functionality. Our purpose in this section is to learn a generative model $g_{\phi}(\widehat{pos}_{T}|c)$, where ${\phi}$ represents the model parameters. This model generates trajectory goals $\widehat{pos}_{T}$ that conform to the distribution of the conditional variable $c$. Seq-CVAE architecture is illustrated in Figure 5.
CVAE was first introduced in the field of image processing [26]. It compresses high-dimensional data, such as pixel-based images, into a low-dimensional latent space while preserving key features of the image. However, unlike images, trajectories typically consists of low-dimensional data, such as the coordinates in $pos_{0:\tau}$. To extract latent features like velocity, direction, and acceleration, these inputs require dimensionality expansion rather than compression. Consequently, we propose modifying the CVAE Encoder’s dimensionality reduction mechanism to perform dimensionality expansion instead.
Furthermore, regarding the construction of condition $c$ in CVAE, Yue et al. [43] extracts features from observed trajectory using an MLP. To better preserve the temporal feature of the trajectory, we adopt an LSTM module. To maintain information balance within the Encoder, the last step hidden state $h_{\tau}$ of the LSTM is passed through a fully connected layer for dimensionality reduction before concatenation, as shown in Figure 5. Therefore, for the Seq-CVAE, the condition $c$ for encoder and decoder is defined as
@@ 4.4 Seq-CVAE: LSTM-Based CVAE
where $FC$ represents the fully connected layer, $h_{\tau}$ denotes the hidden state of LSTM at the last observed time step $\tau$, and $pos_{0:\tau}$ represents the coordinates of observed trajecotry.
@@
5 Experiment
Datasets. We use the well-established datasets in the field of pedestrian trajectory prediction: ETH [22] and UCY [16]. The ETH/UCY datasets consist of five sub-datasets, covering four different scenarios and 1536 trajectories. These pedestrian trajectories include various interaction scenarios such as walking together, as couples, and in groups. Similar to previous work [31, 43, 42], we adopt the leave-one-out strategy for splitting the training and testing sets. That is, four datasets are used for training and validation, and one dataset is used for testing. Since partitioning the neighboring regions requires real-world distances, we convert the coordinate scale to real-world metrics in meters. Due to the different frame rates of the ETH and UCY datasets, we unify both to 2.5 fps, corresponding to a time step of 0.4 seconds.
Evaluation Protocol. Like many previous work [40, 41, 42], visual information is not used in our approach. During the test phase, our model takes the trajectory within 3.2 seconds as input and outputs the trajectory for the subsequent 4.8 seconds. For fair comparison, we adopt the best-of-$K$ protocol, which is sampling $K=20$ times and choose the best result for metric computation. This is the standard protocol in pedestrian trajectory prediction [17, 43, 7, 41].
Metrics. We use Average Displacement Error (ADE) and Final Displacement Error (FDE) as evaluation metrics. ADE refers to the average displacement error of the predicted trajectory $\widehat{pos}_{\tau+1:T}$ compared to the ground truth trajectory $pos_{\tau+1:T}$ and FDE represents the displacement error of the predicted goal $\widehat{pos}_{T}$ compared to the ground truth goal $pos_{T}$.
$FDE=\left\|\widehat{pos}_{T}-pos_{T}\right\|$ (7)
Additionally, to validate the effectiveness of the SSOS strategy, we construct the Initial Displacement Error (IDE) metric, which will be discussed in detail in Section 5.2.
Implementation Details. In our model, the Transformer component adopts the original Transformer architecture [29]. For the Neighbor Encoder module, the hidden size is set to 128. In the Seq-CVAE module, the LSTM hidden size and the latent size are configured to 256, and the MLP parameter design is illustrated in Figure 5. In the data preprocessing stage, the neighbor radius $r$ for social interaction is set to 2 meters, which is determined based on the average ADE performance. Due to significant variations in the walking speed and direction across different scenes, this study employs data augmentation by rotatating trajectories at 0°, 90°, 180°, and 270°, along with scaling the trajectory length by a factor of 2. Notably, the direction-based interaction information $\mathcal{N}$ is transformed accordingly during rotation. To ensure the stability of the data distribution, all trajectories are shifted so that they have start at coordinate (0,0). For the distance features in social interaction information, the reciprocal of the distance is taken, meaning the diminishing influence of pedestrians farther away. These values are subsequently normalized using Min-Max normalization.
During training, the main components of InSyn (Interaction Encoder and Trajectory Generator) and the Seq-CVAE Goal Sampler are trained independently and the teacher forcing technique is applied to facilitate training. For the Encoder-Generator, the SSOS strategy is applied. The learning rate is set to $1\times 10^{-4}$. For the Seq-CVAE, the KL divergence weight in the loss function is configured to 5 to balance the model’s reconstruction capability and latent space distribution requirements, and the learning rate is set to $1\times 10^{-3}$. The model is optimized using the Adam optimizer with 50 epochs. Mean Squared Error (MSE) serves as the loss function. During testing, the model generates predictions autoregressively. All experiments are conducted on a single NVIDIA RTX 4090 GPU, hosted on the AutoDL cloud platform.
5.1 Comparison
Table 1: Results comparison on all datasets. Evaluation is performed using the Best-of-20 protocol strategy. The lower the better. Bold Represents the Best.
@@ 5.1 Comparison
5.1.1 Comparison Baselines
@@
We compare our model with representative methods PECNet [7] and TF [9], as well as the latest work [33, 32, 25, 19]. Additionally, we include some common baselines such as SoPhie [24] and CGNS [17]. The baseline results are from officially-reported metrics.
@@
Among these, SoPhie [24], CGNS [17], STG-DAT [18], and Goal-CurveNet [33] leverage visual information, which our approach does not utilize. TF [9] is the first to apply Transformers for human trajectory prediction; however, unlike our method, it focuses on individual motion without explicitly modeling social or environmental interactions.
5.1.2 Evaluation
@@
The comparative results are demonstrated in Table  1. Compared to the aforementioned methods, our model achieves the best average ADE of 0.26. For goal prediction, the Seq-CVAE achieves excellent performance with an average FDE of 0.52, slightly trailing only behind PECNet [7] and TDAGCN [32] despite its relatively lightweight design. In contrast to FDE, which only evaluates the final predicted position, ADE measures error across all predicted trajectory points, with greater emphasis on intermediate trajectory deviations, thereby better reflecting the model’s capacity to model complex pedestrian interactions. To focus on the model’s ability to capture interactions, we adopt ADE as the evaluation metric for assessing and comparing performance across different scenarios.
@@
Sparse Scenarios. However, on the Hotel dataset, our model underperforms relatively compared to original Transformer [9]. The Hotel dataset is a smaller scene area, which limits the effectiveness of InSyn’s distance-based interaction region partitioning. Moreover, the dataset contains predominantly linear trajectories with sparse interactions. In this scenario, the basic attention mechanism of the original Transformer is sufficient to capture motion patterns effectively, and the additional complex interaction modeling in our model may lead to overfitting, causing excessive reliance on interaction behaviors. This issue is also shared by other complex models such as [32] and [33], as shown in Table 1.
@@
5.2 Ablation Studies
5.2.1 Interaction-Related Components Analysis
For the interaction part, region partition divides the circular area centered on the agent’s position into four regions to capture directional interaction effects. The interaction state, proposed in this study, categorizes interactions into three types: In Sync, Conflict, and No Neighbor. To further investigate the roles of these components in the InSyn model, we deconstruct the model into four variations and evaluate their performance, as summarized in Table 2. These variations allow us to isolate the individual and combined effects of region partition and interaction state. The original InSyn incorporates both region partition and interaction state. Case studies on the variations are provided in the supplementary material.
From the results in Table  2, it is evident that the model incorporating both region rartition and interaction state, namely the original InSyn, achieves the best performance. This demonstrates that explicitly dividing interaction behaviors enhances the model’s ability to extract complex interaction patterns among an agent’s neighbors, leading to more reasonable trajectory predictions.
Specifically, in the absence of region partition, the model may struggle to discern the direction of influence, relying solely on training data distributions to infer future trajectory deviations—a limitation that undermines its generalizability. Similarly, if without interaction state, the model would neglect specific interaction relationships between neighboring pedestrians and judge influence solely based on distance. This would underperform when the training and testing sets differ. For instance, if the training set contains more In Sync scenarios, the trained model may underestimate repulsive effects in Conflict situations.
Table 2: ADE Result of Ablation Study on Region Partition and Interaction State. RP: region partition; IS: interaction state. w/o denotes Without. InSyn incorporates both region partition and interaction state, whereas the Baseline model includes neither.
@@ 5.2.1 Interaction-Related Components Analysis
5.2.2 SSOS Strategy Analysis
Additionally, we conducted an ablation study on the SSOS strategy proposed in this paper. SSOS is designed to mitigate the divergence at the first step when using Transformer’s Encoder-Decoder architecture for numerical time-series prediction tasks, such as pedestrian trajectory prediction. In this context, steps $0:\tau$ represent observed input to the encoder, while time steps $\tau+1:T$ correspond to the future trajectory to be predicted. We replace the traditional single SOS with the sequence-based SSOS. This approach incorporates the observed values from steps $0:\tau$ as inputs for both the encoder and decoder during the initial stage. During the autoregressive process, the outputs are progressively concatenated and fed back into the Trajectory Generator as inputs. This allows the model to predict the complete trajectory while computing the loss for all outputs (Equation 3). To evaluate the effectiveness of SSOS, we introduce the Initial Displacement Error (IDE):
@@ 5.2.2 SSOS Strategy Analysis
Table  3 compares the average IDE results across all datasets using SSOS and SOS (set to $pos_{\tau}$). The results show that the SSOS strategy reduces the average IDE by approximately 6.58% compared to the SOS strategy. This reduction alleviates prediction divergence at time $\tau+1$, enabling smoother transitions between observed and predicted trajectories. Additionally, it mitigates error accumulation during autoregressive prediction.
Table 3: IDE Result of Ablation Study on SSOS. InSyn-SSOS: trained using SSOS; InSyn-SOS: trained using SOS. InSyn-SSOS represents the original InSyn model.
@@
5.3 Visualization
To further explore its practical performance, we conducted a qualitative analysis of the model’s behavior in real-world scenarios via visualization. Figure 6 illustrates the walking state of a specific agent (pedestrian 1) during the observation time steps, with details on its trajectory and interactions. After feeding this input into InSyn, the model outputs the predicted trajectory distribution, which is visualized using Kernel Density Estimation (KDE).
In this example, pedestrian 2 moves in the same direction as the agent, suggesting synchronization. With the exception of a brief Conflict at $t=3$, when pedestrian 2 temporarily moves to the upper-right of the agent, pedestrian 2 remains largely In Sync with the agent. Consequently, despite consistently staying to the agent’s vicinity, pedestrian 2 exerts relatively weak influence on the agent’s trajectory. At $t=7$, pedestrian 3 enters the agent’s neighbor region while moving in the opposite direction. The model encodes this information—"a Conflict from the upper-left region at $t=7$, at a distance of $d_{t=7}^{lu}$"—into a representation. The LSTM aggregates representations across history $0:\tau$ ($\tau=7$) and the Trajectory Generator interprets its hidden state $h_{\tau}$ as an "avoidance" signal. As a result, during the initial prediction steps, the agent turns to the right to avoid pedestrian 3. Since the agent consistently walks toward the upper-right during the observation period, the goals sampled by Seq-CVAE are predominantly directed toward the upper-right. After predicting the "avoidance" behavior, the goal-driven attention mechanism guides the model to refocus on moving toward the sampled goal.
Due to the inherent randomness in pedestrian trajectories (e.g., variations in speed), long-term goal predictions may occasionally deviate from the ground truth. However, intermediate trajectory predictions, including directional changes, remain relatively accurate (see Figure 6). Through visualizing pedestrian trajectories, the model demonstrates its sensitivity to interaction relationships and its ability to effectively leverage goal-driven predictions. More case studies are provided in the supplementary material.
Figure 6: Visualization of Predicted Trajectory Distributions. We visualized the trajectory prediction of an agent (pedestrian 1) in a scene where In Sync, Conflict, and No Neighbor occur. The number of samples is $K$=20, meaning 20 goals were sampled. For each goal, a trajectory was generated in a goal-driven manner. Since pedestrian 4 remains outside the agent’s neighbor region, the model disregards its influence.
6 Conclusion
This paper introduces InSyn, a novel model that extends the Transformer architecture to explicitly capture complex pedestrian interaction patterns. Experimental results demonstrate that our model excels in scenarios involving intricate interaction behaviors, effectively modeling detailed temporal and spatial relationships. Furthermore, to mitigate initial prediction divergence when applying the Transformer’s encoder-decoder framework to numerical sequence data, we propose the SSOS strategy, which relatively alleviates this challenge.

### v2

@@ IV-B SSOS: Seq-Start of Seq
IV-C Seq-CVAE: LSTM-Based CVAE
RNN (typically LSTM [20]) and Conditional Variational Autoencoder (CVAE) [36] have been widely applied in pedestrian trajectory prediction [37, 21, 6, 3]. We combine the two approaches and propose a generator, Seq-CVAE (Sequence CVAE), which is designed for sampling trajectory goals. The model predicts goals given the condition $c$, which is derived from the observed trajectory $pos_{0:\tau}$. These goals are subsequently fed into the Interaction Encoder to achieve its goal-driven functionality. Our purpose in this section is to learn a generative model $g_{\phi}(\widehat{pos}_{T}|c)$, where $\phi$} represents the model parameters. This model generates trajectory goals $\widehat{pos}_{T}$ that conform to the distribution of the conditional variable $c$. The Seq-CVAE architecture is illustrated in Figure 4.
CVAE was first introduced in the field of image processing [36]. It compresses high-dimensional data, such as pixel-based images, into a low-dimensional latent space while preserving key features of the image. However, unlike images, trajectories typically consist of low-dimensional data, such as the coordinates in $pos_{0:\tau}$. To extract latent features like velocity, direction, and acceleration, these inputs require dimensionality expansion rather than compression. Consequently, we propose modifying the CVAE Encoder’s dimensionality reduction mechanism to perform dimensionality expansion instead.
Furthermore, regarding the construction of condition $c$ in CVAE, Yue et al. [6] extract features from observed trajectory using an MLP. To better preserve the temporal feature of the trajectory, we adopt an LSTM module. To maintain information balance within the Encoder, the last step hidden state $h_{\tau}$ of the LSTM is passed through a fully connected layer for dimensionality reduction before concatenation, as shown in Figure 4. Therefore, for the Seq-CVAE, the condition $c$ for the encoder and decoder is defined as
@@ IV-C Seq-CVAE: LSTM-Based CVAE
where $FC$ represents the fully connected layer, $h_{\tau}$ denotes the hidden state of LSTM at the last observed time step $\tau$, and $pos_{0:\tau}$ represents the coordinates of the observed trajectory.
@@
V Experiment
TABLE I: Results comparison on all datasets. Bold: Best
@@ V Experiment
Datasets. We use the well-established datasets in the field of pedestrian trajectory prediction: ETH [12] and UCY [13]. The datasets include five sub-datasets, covering four different scenarios and 1536 trajectories. These trajectories include various interaction scenarios such as walking together, as couples, and in groups. Similar to previous work [43, 6, 34], we adopt the leave-one-out strategy for splitting the training and testing sets. That is, four datasets are used for training and validation, and one dataset is used for testing. Since partitioning the neighboring regions requires real-world distances, we convert the coordinate scale to real-world metrics in meters. Due to the different frame rates of the ETH and UCY datasets, we unify both to 2.5 fps, corresponding to a time step of 0.4 seconds.
Evaluation Protocol. Like many previous work [21, 8, 34], visual information is not used in our approach. During the test phase, our model takes the trajectory within 3.2 seconds as input and outputs the trajectory for the subsequent 4.8 seconds. For fair comparison, we adopt the best-of-$K$ protocol, which is sampling $K=20$ times and choosing the best result for metric computation. This is the standard protocol in pedestrian trajectory prediction [40, 6, 7, 8].
Metrics. We use Average Displacement Error (ADE) and Final Displacement Error (FDE) as evaluation metrics. ADE refers to the average displacement error of the predicted trajectory $\widehat{pos}_{\tau+1:T}$ compared to the ground truth trajectory $pos_{\tau+1:T}$, and FDE represents the displacement error of the predicted goal $\widehat{pos}_{T}$ compared to the ground truth goal $pos_{T}$.
$FDE=\left\|\widehat{pos}_{T}-pos_{T}\right\|$ (7)
Additionally, to validate the effectiveness of the SSOS strategy, we construct the Initial Displacement Error (IDE) metric, which will be discussed in detail in Section V-B.
Implementation Details. In our model, the Transformer component adopts the original Transformer architecture [35]. For the Neighbor Encoder module, the hidden size is set to 128. In the Seq-CVAE module, the LSTM hidden size and the latent size are configured to 256, and the MLP parameter design is illustrated in Figure 4. In the data preprocessing stage, the neighbor radius $r$ for social interaction is set to 2 meters, which is determined based on the average ADE performance. Due to significant variations in the walking speed and direction across different scenes, this study employs data augmentation by rotating trajectories at 0°, 90°, 180°, and 270°, along with scaling the trajectory length by a factor of 2. Notably, the direction-based interaction information $\mathcal{N}$ is transformed accordingly during rotation. To ensure the stability of the data distribution, all trajectories are shifted so that they start at coordinate (0,0). For the distance features in social interaction information, the reciprocal of the distance is taken, meaning the diminishing influence of pedestrians farther away. These values are subsequently normalized using Min-Max normalization.
During training, the main components of InSyn (Interaction Encoder and Trajectory Generator) and the Seq-CVAE Goal Sampler are trained independently, and the teacher forcing technique is applied to facilitate training. For the Encoder-Generator, the SSOS strategy is applied. The learning rate is set to $1\times 10^{-4}$. For the Seq-CVAE, the KL divergence weight in the loss function is configured to 5 to balance the model’s reconstruction capability and latent space distribution requirements, and the learning rate is set to $1\times 10^{-3}$. The model is optimized using the Adam optimizer with 50 epochs. Mean Squared Error (MSE) serves as the loss function. During testing, the model generates predictions autoregressively. All experiments are conducted on a single NVIDIA RTX 4090 GPU, hosted on the AutoDL cloud platform.
V-A Comparison
V-A1 Comparison Baselines
@@
We compare our model with representative methods PECNet [7] and TF [11], as well as the latest work [42, 4, 5, 9]. Additionally, we include some common baselines such as SoPhie [39] and CGNS [40]. The baseline results are from officially reported metrics.
@@
Among these, SoPhie [39], CGNS [40], STG-DAT [27], and Goal-CurveNet [42] leverage visual information, which our approach does not utilize. TF [11] is the first to apply Transformers for human trajectory prediction; however, unlike our method, it focuses on individual motion without considering social interactions.
V-A2 Evaluation
@@
The comparative results are demonstrated in Table I. Compared to the aforementioned methods, our model achieves the best average ADE of 0.26. For goal prediction, the Seq-CVAE achieves excellent performance with an average FDE of 0.52, slightly trailing only behind PECNet [7] and TDAGCN [4] despite its relatively lightweight design. In contrast to FDE, which only evaluates the final predicted position, ADE measures error across all predicted trajectory points, with greater emphasis on intermediate trajectory deviations, thereby better reflecting the model’s capacity to model complex pedestrian interactions. To focus on the model’s ability to capture interactions, we adopt ADE as the evaluation metric for assessing and comparing performance across different scenarios.
@@
Sparse Scenarios. However, on the Hotel dataset, our model underperforms relatively compared to the original Transformer [11]. The Hotel dataset is a smaller scene area, which limits the effectiveness of InSyn’s distance-based interaction region partitioning. Moreover, the dataset contains predominantly linear trajectories with sparse interactions. In this scenario, the basic attention mechanism of the original Transformer is sufficient to capture motion patterns effectively, and the additional complex interaction modeling in our model may lead to overfitting, causing excessive reliance on interaction behaviors. This issue is also shared by other complex models such as [4] and [42], as shown in Table I.
@@
V-B Ablation Studies
V-B1 Interaction-Related Components Analysis
For the interaction part, the region partition divides the circular area centered on the agent’s position into four regions to capture directional interaction effects. The interaction state, proposed in this study, categorizes interactions into three types: In Sync, Conflict, and No Neighbor. To further investigate the roles of these components in the InSyn model, we deconstruct the model into four variations and evaluate their performance, as summarized in Table II. These variations allow us to isolate the individual and combined effects of region partition and interaction state. The original InSyn incorporates both region partition and interaction state. Case studies on the variations are provided in the supplementary material.
From the results in Table  II, it is evident that the model incorporating both region partition and interaction state, namely the original InSyn, achieves the best performance. This demonstrates that explicitly dividing interaction behaviors enhances the model’s ability to extract complex interaction patterns among an agent’s neighbors, leading to more reasonable trajectory predictions.
Specifically, in the absence of region partition, the model may struggle to discern the direction of influence, relying solely on training data distributions to infer future trajectory deviations—a limitation that undermines its generalizability. Similarly, without interaction state, the model would neglect specific interaction relationships between neighboring pedestrians and judge influence solely based on distance. This would underperform when the training and testing sets differ. For instance, if the training set contains more In Sync scenarios, the trained model may underestimate repulsive effects in Conflict situations.
TABLE II: ADE Result of Ablation Study on Region Partition (RP) and Interaction State (IS)
@@ V-B1 Interaction-Related Components Analysis
V-B2 SSOS Strategy Analysis
Additionally, we conducted an ablation study on the SSOS strategy proposed in this paper. SSOS is designed to mitigate the divergence at the first step when using the Transformer’s Encoder-Decoder architecture for numerical time series prediction tasks, such as pedestrian trajectory prediction. In this context, steps $0:\tau$ represent observed input to the encoder, while time steps $\tau+1:T$ correspond to the future trajectory to be predicted. We replace the traditional single SOS with the sequence-based SSOS. This approach incorporates the observed values from steps $0:\tau$ as inputs for both the encoder and decoder during the initial stage. During the autoregressive process, the outputs are progressively concatenated and fed back into the Trajectory Generator as inputs. This allows the model to predict the complete trajectory while computing the loss for all outputs (Equation 3). To evaluate the effectiveness of SSOS, we introduce the Initial Displacement Error (IDE):
@@ V-B2 SSOS Strategy Analysis
Table III shows the average ADE results across all datasets, comparing different ablations of our model. ”w/o” refers to ”without” a component. The full model (InSyn) includes both RP and IS, while the Baseline model excludes both components. The results show that the SSOS strategy reduces the average IDE by approximately 6.58% compared to the SOS strategy. This reduction alleviates prediction divergence at time $\tau+1$, enabling smoother transitions between observed and predicted trajectories. Additionally, it mitigates error accumulation during autoregressive prediction.
TABLE III: IDE Result of Ablation Study on SSOS
@@
V-C Case Study
Figure 5: Case Study of Region Partition and Interaction State. This figure compares three variants of InSyn: (1) without region partition (w/o-RP), (2) without interaction state (w/o-IS), and (3) the full model incorporating both components (InSyn). To isolate the effect of interaction modeling, we exclude the Goal Sampler in this evaluation.
To further investigate the effectiveness of region partition and interaction state, we conduct a Case study. The cases (see Figure 5) cover various interaction scenarios, including paired walking, group walking, and pedestrian conflict.
In Sync Scenarios. Cases (a), (b), (e), and (f) represent instances where pedestrians walk in coordination with others nearby. In cases (a) and (b), where the target pedestrian moves alongside a companion, InSyn preserves the direction and velocity of the agent’s motion at the end of the observation period, closely matching the ground truth. In contrast, both ablated variants produce noticeable deviations. These cases suggest that the influence of the neighboring pedestrian is relatively weak. InSyn effectively models the behaviors by accurately capturing the reduced impact of the neighbor on the agent’s motion. Case (e) depicts a group walking situation with multiple neighboring pedestrians moving in sync. Here, only InSyn predicts the correct direction of motion, while both variants mistakenly infer an opposite trajectory, highlighting their failure to capture the low-impact nature of synchronized group behavior. Case (f) presents a simple, slow-moving scenario where all models achieve comparable performance, showing that the base model alone suffices for simple prediction tasks.
Conflict Scenarios. Cases (c) and (d) illustrate interactions where conflict plays a dominant role. In case (c), although the agent’s last observed motion trends slightly upward, InSyn predicts a flatter, rightward trajectory that better aligns with the ground truth. This adjustment reflects the model’s ability to recognize a conflicting pedestrian approaching from above, thereby reducing unnecessary upward movement. The w/o-RP and w/o-IS variants, lacking directional awareness and interaction-type recognition respectively, simply extrapolate the upward motion and thus produce larger errors. Case (d), however, reveals a limitation of InSyn. In this instance, the model appears to overestimate the influence of a pedestrian approaching from below, leading to an excessive upward deviation from the true path. This suggests that while InSyn improves reliability and interpretability in most complex scenarios, further refinement is needed to enhance control precision in handling conflict interactions.
These case studies demonstrate that modeling explicit interaction types and directional influence contributes to more trustworthy and socially-aware predictions. Compared to black-box representations, our approach enhances interpretability, allowing for more transparent reasoning about how different pedestrian behaviors shape trajectory outcomes. This interpretability not only builds trust in the model’s decisions but also provides actionable insights for refining interaction modeling strategies.
VI Conclusion
This paper presents InSyn, a trustworthy trajectory prediction model that extends the Transformer architecture to explicitly capture pedestrian complex interaction patterns. By introducing interpretable structure into otherwise black-box models, InSyn improves interpretability and robustness in socially complex environments. Experimental results validate the effectiveness of our approach in both dense and sparse scenarios, offering more reliable predictions. Besides, our proposed SSOS strategy enhances prediction accuracy by mitigating initial-step divergence in numerical sequence modeling. A detailed case study further illustrates how modeling explicit interaction patterns contributes to interpretable and trustworthy predictions in diverse real-world scenarios.